A 4.5 m long rod has a mass of 2.8 kg and is stretched 0.21 m by force. What is the strain on the rod?

Select one:

a) 21.43     b) 0.047     c) 0.13    d) 6.1

Assuming atmospheric pressure to be 1.01 x 105 Pa and the density of sea water to be 1025 kg/m3 , What is the absolute pressure at a depth of 15.0 m below the surface of the ocean?

Select one:

a. 1.508 x 105 Pa

b. 1.16 x 105 Pa

c. 4.98 x 104 Pa

d. 2.52 x 105 Pa
Answer #1

Answer: (i)

Length of the rod L=4.5~m

Change in lenght, \Delta L=0.21~m

Strain, \epsilon =\frac{\Delta L}{L}=\frac{0.21}{4.5}=0.047

(ii) Atmospheric pressure, P_A=1.01\times10^5~Pa

density of the water, \rho =1025~kg/m^3

depth, h=15.0~m

absolute pressure at a depth h below the surface of the ocean,

P=P_A+\rho gh=(1.01\times10^5~Pa)+(1025\times9.8\times15.0)~Pa

=(1.01\times10^5~Pa)+(1.51\times10^5~Pa)=2.52\times10^5~Pa